# Theory of Irreversibility in Quantum Many-Body Systems

### v2

@@ I Introduction
In classical systems, the probability density of an ensemble of trajectories evolves in time under the action of the Frobenius-Perron (FP) operator. Because of the conservation of phase-space volume (Liouville’s theorem), the evolution is unitary and the eigenvalues of the FP operator lie on the unit circle in the complex plane. However, if the system is chaotic, the spectrum of the FP operator becomes dense, and isolated singularities can emerge inside the unit circle upon analytic continuation, the so-called Ruelle-Pollicott (RP) resonances [4, 5, 6, 7]. These resonances control the decay of different observables, with the associated timescales inversely proportional to their distance to the unit circle, with the leading RP resonance ruling the late-time approach to a steady state. Thus, the singularities arising from chaos lead to irreversibility in purely unitary dynamics [2, 3].
@@
Despite all this recent progress, the theory of quantum many-body RP resonances is still rudimentary. The existence of quantum RP resonances can be rigorously established in certain single-particle systems [31, 9, 10, 11, 12], but most of the scarce evidence in interacting many-body systems is heuristic or numerical. The notable exception is the dissipative random phase model (DRPM), a quantum-chaotic dissipative quantum circuit model introduced by the same authors of the present paper in Ref. [17]. In the limit of large local Hilbert space dimension, this model is exactly solvable for arbitrary system size and dissipation strength and constitutes, therefore, a perfect testbed to further develop the theory of irreversibility of many-body quantum systems. In this paper, we proceed in this direction with four distinct goals in mind.
@@
Third, we study a so-far unexplored dynamical regime of many-body quantum chaos, the Ruelle regime. Arguably, the two most striking features of chaos are sensitivity to initial conditions and mixing. If a quantum system has a well-defined semiclassical regime, the former manifests as an early-time exponential growth in out-of-time-order correlators (OTOCs), which defines a quantum Lyapunov exponent. The Lyapunov regime has been extensively investigated for both single-particle [33, 34, 35, 12] and many-body systems [36, 37], but many systems without an effective semiclassical parameter completely lack this behavior. On the contrary, the decay of correlation functions is a universal feature of all chaotic quantum systems—the quantum version of mixing. To illustrate this point, we analytically demonstrate that, while the OTOC of the RPM has no Lyapunov regime, the same quantum RP resonances as before rule the saturation of the OTOC [38] (and its decay in the presence of dissipation), which is identified as the Ruelle regime of the model.
@@ II.2 Quantum Ruelle-Pollicott resonances
We order the eigenvalues as $0\leq|\mathrm{Re}\,\nu_{0}(\gamma)|\leq\cdots\leq|\mathrm{Re}\,\nu_{D^{2}-1}(\gamma)|$. In terms of the basis operators, any observable $\mathcal{O}$ can be then decomposed as $|\mathcal{O}\rangle\!\rangle=\sum_{k}c_{k}|R_{k}\rangle\!\rangle$ where $c_{k}=\langle\!\langle L|\mathcal{O}\rangle\!\rangle={\rm Tr}(L^{\dagger}_{k}\mathcal{O})$, which implies that
$\mathcal{O}(t)=(\mathcal{W}^{\dagger})^{t}[\mathcal{O}(0)]=\sum_{k}c_{k}e^{-\nu_{k}(\gamma)t}|R_{k}\rangle\!\rangle.$ (11)
@@
$\langle\mathcal{O}(t)\mathcal{O}(0)\rangle\sim e^{-\nu_{1}t}.$ (13)
@@ II.3 Irreversibility in Floquet many-body quantum systems
II.3.1 Ensemble-averaged autocorrelation functions
@@
In this paper, we shall be mainly concerned with the temporal behavior of the ensemble-averaged autocorrelation functions of basis operators $P_{\alpha}$,
$\overline{C_{\alpha\alpha}(t)}=\overline{\langle P_{\alpha}(t)P_{\alpha}\rangle},$ (14)
where $P_{\alpha}(t)=(\mathcal{W}^{\dagger})^{t}[P_{\alpha}]$ is the Heisenberg evolution of $P_{\alpha}$. Note that Eq. (14) is valid for both open and closed systems; in the latter case, it reduces to $P_{\alpha}(t)=(W^{\dagger})^{t}P_{\alpha}W^{t}$.
@@ II.3.3 Autocorrelation-function expansion of the SFF
First, they decay from their initial value $\overline{C_{\alpha\alpha}(0)}=1$ to a local minimum after a characteristic decay timescale $t_{\mathrm{d},\alpha}$ of the order of the Floquet period, which follows from the cancelation of system-dependent fluctuations upon ensemble-averaging. Second, the decay does not extend indefinitely and $\overline{C_{\alpha\alpha}(t)}$ grows sharply after $t_{\mathrm{d},\alpha}$, before decaying again. This peak, which we call the Thouless peak [17], is due to interactions of neighboring sites and is an exclusive signature of many-body physics. Third, after the Thouless peak decays to a value $\overline{C_{\alpha\alpha}}\ll q^{-L}$, the autocorrelation functions start regrowing linearly after a ramp timescale $t_{\mathrm{r},\alpha}$, before eventually plateauing at $\overline{C_{\alpha\alpha}(t>t_{\mathrm{H}})}=q^{-L}$ at the Heisenberg time $t_{\mathrm{H}}=q^{L}$.
@@
From the previous discussion, it becomes clear that to obtain RP resonances (i.e., the decay rates of individual correlation functions) from eigenvalues of the evolution operator (i.e., the late-time behavior of the SFF) we must suppress the ramp of the SFF. It is, therefore, crucial to identify which operators contribute to the ramp of the SFF in the thermodynamic limit and devise a scheme to suppress their growth.
@@ II.4 Models
Based on the structure of the RPM, in the DRPM each Floquet step is accompanied by a quantum channel that introduces dissipation. For simplicity, we assume that the quantum channel has a site-resolved form $\Phi=\bigotimes_{x}\Phi_{x}$. The action of the single-site channel $\Phi_{x}$ on the state $\varrho_{x}$ at site $x$ can be expressed in terms of the Kraus decomposition $\Phi_{x}(\varrho_{x})=\sum_{i=0}^{k-1}M_{i}\varrho_{x}M^{\dagger}_{i}$, where $k$ is the number of channels, with Kraus operators $M_{i}$ that are normalized as $q^{-1}{\rm Tr}(M_{i}M^{\dagger}_{j})=\eta_{i}\delta_{ij}$ with some channel-dependent $10<\eta_{i}<$ 1. They satisfy the sum rule $\sum_{i=0}^{k-1}M^{\dagger}_{i}M_{i}=I_{q}$ where $I_{q}$ is the $q\times q$ identity matrix. For example, the depolarizing channel with depolarizing probability $p$, $\Phi_{x}(\varrho_{x})=(1-p)\varrho_{x}+pI_{q}/q$, has $\eta_{0}=1-p(q^{2}-1)/q^{2}$ ($\eta_{0}\approx 1-p$ when $q\to\infty$) and $\eta_{i}=p/q^{2}$ [17]. It is then convenient to define an effective dissipation strength $\mathbb{R}^{+}\ni\gamma=-\varepsilon^{-1}\log\eta_{M}$, where $\eta_{M}$ is the largest of the $\eta_{i}$. $\gamma=0$ corresponds to the dissipationless limit and $\gamma\to\infty$ to infinitely-strong dissipation. By slight abuse of notation, let us also denote the quantum channel acting on vectorized states by $\Phi$. The whole time evolution of the circuit can then be represented by $\mathcal{W}^{t}$ where $\mathcal{W}=\Phi(W\otimes W^{*})$.
@@ II.5 Analytic results
where $\mathcal{O}(y,t)$ is the time evolution of a local traceless operator $\mathcal{O}$, which, at $t=0$, is supported at position $y$. In closed systems, like the RPM, the norm $\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle$ is a constant (for simplicity, we normalize $\mathcal{O}$ such that $\langle\mathcal{O}^{2}\rangle=1$) to which the OTOC relaxes. The relaxation rate is thus determined by the decay of the four-point contribution $\langle\mathcal{O}(y,t)P_{\alpha^{x}}\mathcal{O}(y,t)P_{\alpha^{x}}\rangle$. We find that the OTOC shows two-stage relaxation [73, 74, 75, 76, 77, 78, 79, 80, 81] for PBCs but it does not for OBCs. Namely, when PBCs are imposed, the relaxation rate during the first stage, which takes place over the timescale $t\lesssim L/v_{B}$ where $v_{B}$ is the butterfly velocity, is given by $2\varepsilon$, whereas in the second stage, for $t\gtrsim L/v_{B}$, it increases to $4\varepsilon$. The large-$q$ OTOC during the second stage thus reads asymptotically
@@ III.1 Autocorrelation functions
that acts on the $t$-dimensional vector space spanned by the cyclic pairing states $|s\rangle$. Here, $\mathtt{E}$ is the constant matrix of ones. We can thus replace the (1+1)-dimensional evolution of the physical degrees of freedom with $q$-dimensional local Hilbert space by a one-dimensional evolution of pairing degrees of freedom with $t$-dimensional local Hilbert space; these pairing degrees of freedom evolve in space under the action of a transfer matrix $T$. Because of the Gaussian phase interaction between neighboring sites, the transfer matrix acts in a very simple way: if the two adjacent sites have the same pairing, then the transfer matrix does nothing; if the pairings are different—i.e., there is a pairing domain wall—, they incur an entropic cost $e^{-\varepsilon t}$. It is this cost of domain walls that causes the decay of correlation functions in the RPM.
@@ III.3 Operator truncation and the truncated form factor
Previously we claimed that the RP resonances obtained in the weak-dissipation limit of the DRPM are an intrinsic property of the closed RPM. Then, the value of the RP resonance must be independent of the “seed” of nonunitarity. In this section, we show that this is indeed the case by using a different source of nonunitary, namely, operator truncation (or coarse-graining), as originally proposed by Prosen [14, 15] for the kicked Ising model, see also Ref. [18, 19] for recent studies. For the RPM we perform the computation analytically and find the same value of the RP resonance, thus proving the agreement of the two methods for the first time.
@@ III.4.1 Relaxation of the OTOC in the RPM
We thus infer that during the second stage of relaxation, the decay rate is replaced by $4\varepsilon$, which is twice larger than that during the first stage. Note that this rate can also be obtained by simply invoking the relation between RP resonances and late-time dynamics, as discussed in Sec. II.2. Namely at late times we can effectively assume $|\mathcal{O}(y,t)\rangle\!\rangle\approx e^{-\nu_{1}t}|R_{1}\rangle\!\rangle$. Putting this into Eq. (48), we immediately observe that the ensemble-averaged OTOC decays to one with the rate $2\nu_{1}$, which reads $4\varepsilon$ for the RPM with PBCs. Thus, also the saturation of the OTOC in the large-$q$ RPM is ruled by (twice) the leading RP resonance, which is in agreement with the assertion that the RP resonances are intrinsic properties of the closed system.
@@
$\langle\mathcal{O}$ $(y,t)P_{\alpha^{x}}\mathcal{O}(y,t)P_{\alpha^{x}}\rangle$
@@
$\langle\mathcal{O}(y,t)P_{\alpha^{x}}\mathcal{O}(y,t)P_{\alpha^{x}}\rangle=\widehat{S}^{x-y}_{0t}.$ (59)
This indicates that the OTOC always decays with the same rate $2\varepsilon$ for any system size and the value of $y$, hence the two-stage relaxation is absent with OBCs. Note that, as in the case of PBCs, the rate associated with the long-time decay with finite $L$ can be deduced solely based on the fact that late-time dynamics is controlled by the leading RP resonance, which is $\varepsilon$ for the RPM with OBCs. Following the same argument above for PBCs, the OTOC then has to decay with the rate $2\varepsilon$ (since it has two time-evolved operators $\mathcal{O}(0,t)$), which is precisely the case in Eq. (59).
@@ III.4.2 Decay of the OTOC in the DRPM
Similarly, we can also express the norm $\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle$, which is independent of $y$ for PBCs, as
$\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle=\mathrm{Tr}[(E_{\mathrm{d}}S)^{L}|0\rangle\langle 0|]=[\widehat{S}^{L}_{\mathrm{d}}]_{00}.$ (62)
@@
Due to dissipation, the OTOC $\overline{\mathcal{C}(x,y;t)}$ now decays exponentially to zero instead of approaching unity at long times. As in the DFF, we are particularly interested in taking the large system size limit $L\to\infty$ first and then sending $\gamma\to 0$ (i.e., $\omega\to 1$), under which not only the second term in Eq. (63) but also the first term, which is the norm $\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle$, decays exponentially.
@@
We now show that the two contributions to the OTOC decay with the same rate, which coincides with (twice) the leading RP resonance $2\varepsilon$. This feature of the OTOC was observed before for single-particle systems [12], but our results show that it also extends to the many-body setting, suggesting its universality in quantum dynamics. The transfer-matrix expressions in Eqs. (III.4.1) and (62) are exact but their large-$L$, large-$t$ asymptotics are hard to analyze. Nevertheless, we have already shown in the previous section that, in the dissipationless limit the decay rate of the four-point contribution $\langle\mathcal{O}(y,t)P_{\alpha^{x}}\mathcal{O}(y,t)P_{\alpha^{x}}\rangle$ is indeed given by $2\varepsilon$.
@@
Figure 5: Exact transfer-matrix result for the decay of the norm. The colored lines give the decay rate $\Delta_{\mathrm{norm}}$ of $\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle$ for PBCs in units of $\varepsilon=1$, obtained for $140120<t<$ 140 and $8020<L<$ 80, as a function of $\gamma$. The dashed black line is given by the extrapolation of the collapsed $\Delta_{\mathrm{norm}}(\gamma)$ down to $\gamma\to 0$, where it intercepts the vertical axis at $\Delta_{\mathrm{norm}}(0)=4$.
@@
We thus proceed with the operator norm contribution $\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle$, Eq. (62). Exact evaluation of $[\widehat{S}^{L}_{\mathrm{d}}]_{00}$ shows that for $L>2(t-1)$ it becomes independent of $L$ and we have
$\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle=(\omega\rho^{2})^{(t-1)}\left[\left(1-\rho^{-1};\omega\right)_{t-1}\right]^{2},$ (64)
@@
where $(a;q)_{n}=\prod_{k=0}^{n-1}(1-aq^{k})$ is the $q$-Pochhammer symbol. Notice that here dissipation acts as a $q$-deformation of the closed-system operator norm. Taking the limit $\omega\to 1$ (i.e., $\gamma=0$) reduces it to the result of the previous section, $\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle=1$ regardless of whether the thermodynamic limit is taken before or after, since it is independent of $L$. Instead, we are interested in the long-time decay of the correlator and therefore have to take the limit $t\to\infty$ before $L\to\infty$, which is incompatible with the restriction $L>2(t-1)$ in Eq. (64). In the opposite case, $2(t-1)>L$, we are unable to find a simple closed-form expression. However, we can still evaluate Eq. (62) exactly for arbitrary $t$, $L$, and $\omega$ and extract the decay rate by fitting the late-time decay to an exponential, $\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle=A_{\mathrm{norm}}e^{-\Delta_{\mathrm{norm}}t}$, with $A_{\mathrm{norm}}$ and $\Delta_{\mathrm{norm}}$ fitting parameters, as shown in Fig. 5. We find the behavior expected for anomalous relaxation [17]: the curves for different $L$ collapse to a universal curve up to a critical value of the dissipation strength below which the system sees the finite spacing of the lattice and, correspondingly, the gap closes. As $L$ increases (but keeping it below the times used for the fit), this critical point decreases until reaching $p=0$ in the thermodynamic limit. Extrapolating the collapsed curve to this limit, we obtain a decay rate of $\Delta_{\mathrm{norm}}\approx 4\varepsilon$ for $\gamma=0$, as expected.

### v3

@@ I Introduction
In classical systems, the probability density of an ensemble of trajectories evolves in time under the action of the Frobenius-Perron (FP) operator. Because of the conservation of phase-space volume (Liouville’s theorem), the evolution is unitary and the eigenvalues of the FP operator lie on the unit circle in the complex plane. However, if the system is chaotic, the spectrum of the FP operator becomes dense, and isolated singularities can emerge inside the unit circle upon analytic continuation, the so-called Ruelle-Pollicott (RP) resonances [4, 5, 6, 7]. These resonances control the decay of different observables, with the associated timescales inversely proportional to their distance to the unit circle, and the leading RP resonance rules the late-time approach to a steady state. Thus, the singularities arising from chaos lead to irreversibility in purely unitary dynamics [2, 3].
@@
Despite all this recent progress, the theory of quantum many-body RP resonances is still rudimentary. The existence of quantum RP resonances can be rigorously established in single-particle systems [31, 9, 10, 11, 12] and is well-understood. Notwithstanding, the presence of local interactions in many-body quantum systems renders many-body quantum chaos fundamentally different from its single-particle counterpart. It thus remains unclear if the different manifestations of quantum many-body RP resonances that have recently appeared in the literature are equivalent and how they relate to the different timescales of many-body quantum chaos. While most of the scarce evidence in interacting many-body systems is heuristic or numerical, the dissipative random phase model (DRPM), a quantum-chaotic dissipative quantum circuit model introduced by the same authors of the present paper in Ref. [17], is a notable exception. In the limit of large local Hilbert space dimension, this model is exactly solvable for arbitrary system size and dissipation strength and constitutes, therefore, a perfect testbed to further develop the theory of irreversibility of many-body quantum systems. In this paper, we proceed in this direction with four distinct goals in mind.
@@
Third, we study a so-far unexplored dynamical regime of many-body quantum chaos, the Ruelle regime. Arguably, the two most striking features of chaos are sensitivity to initial conditions and mixing. If a quantum system has a well-defined semiclassical regime, the former manifests as an early-time exponential growth in out-of-time-order correlators (OTOCs), which defines a quantum Lyapunov exponent. The Lyapunov regime has been extensively investigated for both single-particle [33, 34, 35, 12] and many-body systems [36, 37], but many systems without an effective semiclassical parameter completely lack this behavior. On the contrary, the decay of correlation functions is a universal feature of all chaotic quantum systems—the quantum version of mixing. The characterization of chaos at intermediate timescales proportional to the system size—much later than the initial growth of correlations but much earlier than the ultralong timescales characterized by random matrix universaility—has remained essentially unexplored and, crucially, this is the regime where the effect of locality (i.e., many-body interactions) becomes the starkest. To illustrate this point, we analytically demonstrate that, while the OTOC of the RPM has no Lyapunov regime, the same quantum RP resonances as before rule the saturation of the OTOC [38] (and its decay in the presence of dissipation), which is identified as the Ruelle regime of the model.
@@ II.2 Quantum Ruelle-Pollicott resonances
We order the eigenvalues as $0\leq|\mathrm{Re}\,\nu_{0}(\gamma)|\leq\cdots\leq|\mathrm{Re}\,\nu_{D^{2}-1}(\gamma)|$. In terms of the basis operators, any observable $\mathcal{O}$ can be then decomposed as $|\mathcal{O}\rangle\!\rangle=\sum_{k}c_{k}|L_{k}\rangle\!\rangle$ where $c_{k}=\langle\!\langle R_{k}|\mathcal{O}\rangle\!\rangle={\rm Tr}(R^{\dagger}_{k}\mathcal{O})$, which implies that
$\mathcal{O}(t)=(\mathcal{W}^{\dagger})^{t}[\mathcal{O}(0)]=\sum_{k}c_{k}e^{-\nu_{k}^{*}(\gamma)t}|L_{k}\rangle\!\rangle.$ (11)
@@
$\langle\mathcal{O}(t)\mathcal{O}(0)\rangle\sim e^{-\nu_{1}^{*}t}.$ (13)
@@ II.3 Irreversibility in Floquet many-body quantum systems
II.3.1 Autocorrelation functions
@@
In this paper, we shall be mainly concerned with the temporal behavior of the autocorrelation functions of basis operators $P_{\alpha}$,
$C_{\alpha\alpha}(t)=\langle P_{\alpha}(t)P_{\alpha}\rangle,$ (14)
where $P_{\alpha}(t)=(\mathcal{W}^{\dagger})^{t}[P_{\alpha}]$ is the Heisenberg evolution of $P_{\alpha}$ and we recall that $\langle\cdots\rangle=D^{-1}{\rm Tr}(\cdots)$ is the infinite-temperature average. Note that Eq. (14) is valid for both open and closed systems; in the latter case, it reduces to $P_{\alpha}(t)=(W^{\dagger})^{t}P_{\alpha}W^{t}$. We denote the average of the autocorrelation function $C_{\alpha\alpha}(t)$ over an ensemble of similar systems by $\overline{C_{\alpha\alpha}(t)}$.
@@ II.3.3 Autocorrelation-function expansion of the SFF
First, they decay from their initial value $\overline{C_{\alpha\alpha}(0)}=1$ to a local minimum after a characteristic decay timescale $t_{\mathrm{d},\alpha}$ of the order of the Floquet period, which follows from the cancellation of system-dependent fluctuations upon ensemble-averaging. Second, the decay does not extend indefinitely and $\overline{C_{\alpha\alpha}(t)}$ grows sharply after $t_{\mathrm{d},\alpha}$, before decaying again. This peak, which we call the Thouless peak [17], is due to interactions of neighboring sites and is an exclusive signature of many-body physics. Third, after the Thouless peak decays to a value $\overline{C_{\alpha\alpha}}\ll q^{-L}$, the autocorrelation functions start regrowing linearly after a ramp timescale $t_{\mathrm{r},\alpha}$, before eventually plateauing at $\overline{C_{\alpha\alpha}(t>t_{\mathrm{H}})}=q^{-L}$ at the Heisenberg time $t_{\mathrm{H}}=q^{L}$.
@@
From the previous discussion, it becomes clear that to obtain RP resonances (i.e., the decay rates of individual correlation functions) from eigenvalues of the evolution operator (i.e., the late-time behavior of the SFF) we must suppress the ramp of the SFF. It is crucial, therefore, to identify which operators contribute to the ramp of the SFF in the thermodynamic limit and devise a scheme to suppress their growth.
@@ II.4 Models
Based on the structure of the RPM, in the DRPM each Floquet step is accompanied by a quantum channel that introduces dissipation. For simplicity, we assume that the quantum channel has a site-resolved form $\Phi=\bigotimes_{x}\Phi_{x}$. The action of the single-site channel $\Phi_{x}$ on the state $\varrho_{x}$ at site $x$ can be expressed in terms of the Kraus decomposition $\Phi_{x}(\varrho_{x})=\sum_{i=0}^{k-1}M_{i}\varrho_{x}M^{\dagger}_{i}$, where $k$ is the number of channels, with Kraus operators $M_{i}$ that are normalized as $q^{-1}{\rm Tr}(M_{i}M^{\dagger}_{j})=\eta_{i}\delta_{ij}$ with some channel-dependent $10<\eta_{i}<$ 1. They satisfy the sum rule $\sum_{i=0}^{k-1}M^{\dagger}_{i}M_{i}=I_{q}$ where $I_{q}$ is the $q\times q$ identity matrix. For example, the depolarizing channel with depolarizing probability $p$, $\Phi_{x}(\varrho_{x})=(1-p)\varrho_{x}+pI_{q}/q$, has $\eta_{0}=1-p(q^{2}-1)/q^{2}$ ($\eta_{0}\approx 1-p$ when $q\to\infty$) and $\eta_{i}=p/q^{2}$ [17]. It is then convenient to define an effective dissipation strength $\mathbb{R}^{+}\ni\gamma=-\varepsilon^{-1}\log\eta_{M}$, where $\eta_{M}$ is the largest of the $\eta_{i}$. $\gamma=0$ corresponds to the dissipationless limit and $\gamma\to\infty$ to infinitely-strong dissipation. With a slight abuse of notation, let us also denote the quantum channel acting on vectorized states by $\Phi$. The whole time evolution of the circuit can then be represented by $\mathcal{W}^{t}$ where $\mathcal{W}=\Phi(W\otimes W^{*})$.
@@ II.5 Analytic results
where $\mathcal{O}(y,t)$ is the time evolution of a local traceless operator $\mathcal{O}$, which, at $t=0$, is supported at position $y$. In closed systems, like the RPM, the norm $\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle$ is a constant (for simplicity, we normalize $\mathcal{O}$ such that $\langle\mathcal{O}^{2}\rangle=1$) to which the OTOC relaxes. The relaxation rate is thus determined by the decay of the four-point contribution $\langle\mathcal{O}(y,t)P_{\alpha^{x}}\mathcal{O}(y,t)P_{\alpha^{x}}\rangle$. We find that the OTOC shows two-stage relaxation [73, 74, 75, 76, 77, 78, 79, 80, 81] for PBCs but not for OBCs. Namely, when PBCs are imposed, the relaxation rate during the first stage, which takes place over the timescale $t\lesssim L/v_{B}$ where $v_{B}$ is the butterfly velocity, is given by $2\varepsilon$, whereas in the second stage, for $t\gtrsim L/v_{B}$, it increases to $4\varepsilon$. The large-$q$ OTOC during the second stage thus reads asymptotically
@@ III.1 Autocorrelation functions
that acts on the $t$-dimensional vector space spanned by the cyclic pairing states $|s\rangle$. Here, $\mathtt{E}$ is the constant matrix of 1s. We can thus replace the (1+1)-dimensional evolution of the physical degrees of freedom with $q$-dimensional local Hilbert space by a one-dimensional evolution of pairing degrees of freedom with $t$-dimensional local Hilbert space; these pairing degrees of freedom evolve in space under the action of a transfer matrix $T$. Because of the Gaussian phase interaction between neighboring sites, the transfer matrix acts in a very simple way: if the two adjacent sites have the same pairing, then the transfer matrix does nothing; if the pairings are different—i.e., there is a pairing domain wall—, they incur an entropic cost $e^{-\varepsilon t}$. It is this cost of domain walls that causes the decay of correlation functions in the RPM.
@@ III.3 Operator truncation and the truncated form factor
Previously we claimed that the RP resonances obtained in the weak-dissipation limit of the DRPM are an intrinsic property of the closed RPM. Then, the value of the RP resonance must be independent of the “seed” of nonunitarity. In this section, we show that this is indeed the case by using a different source of nonunitary, namely, operator truncation (or coarse-graining), as originally proposed by Prosen [14, 15] for the kicked Ising model; see also Refs. [18, 19] for recent studies. For the RPM we perform the computation analytically and find the same value of the RP resonance, thus proving the agreement of the two methods for the first time.
@@ III.4.1 Relaxation of the OTOC in the RPM
We thus infer that during the second stage of relaxation, the decay rate is replaced by $4\varepsilon$, which is twice as large as that during the first stage. Note that this rate can also be obtained by simply invoking the relation between RP resonances and late-time dynamics, as discussed in Sec. II.2. Namely at late times we can effectively assume $|\mathcal{O}(y,t)\rangle\!\rangle\approx e^{-\nu^{*}_{1}t}|L_{1}\rangle\!\rangle$. Putting this into Eq. (48), we immediately observe that the ensemble-averaged OTOC decays to one with the rate $2\nu^{*}_{1}$, which reads $4\varepsilon$ for the RPM with PBCs. Thus, also the saturation of the OTOC in the large-$q$ RPM is ruled by (twice) the leading RP resonance, which is in agreement with the assertion that the RP resonances are intrinsic properties of the closed system.
@@
$\overline{\langle\mathcal{O}(y,t)P_{\alpha^{x}}\mathcal{O}(y,t)P_{\alpha^{x}}\rangle}=\widehat{S}^{x-y}_{0t}.$ (59)
This indicates that the OTOC always decays with the same rate $2\varepsilon$ for any system size and the value of $y$, hence the two-stage relaxation is absent with OBCs. Note that, as in the case of PBCs, the rate associated with the long-time decay with finite $L$ can be deduced solely based on the fact that late-time dynamics is controlled by the leading RP resonance, which is $\varepsilon$ for the RPM with OBCs. Following the same argument above for PBCs, the OTOC then has to decay with the rate $2\varepsilon$ [since it has two time-evolved operators $\mathcal{O}(0,t)$], which is precisely the case in Eq. (59).
@@ III.4.2 Decay of the OTOC in the DRPM
Similarly, we can also express the norm $\overline{\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle}$, which is independent of $y$ for PBCs, as
$\overline{\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle}=\mathrm{Tr}[(E_{\mathrm{d}}S)^{L}|0\rangle\langle 0|]=[\widehat{S}^{L}_{\mathrm{d}}]_{00}.$ (62)
@@
Due to dissipation, the OTOC $\overline{\mathcal{C}(x,y;t)}$ now decays exponentially to zero instead of approaching unity at long times. As in the DFF, we are particularly interested in taking the large system size limit $L\to\infty$ first and then sending $\gamma\to 0$ (i.e., $\omega\to 1$), under which not only the second term in Eq. (63) but also the first term, which is the norm $\overline{\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle}$, decays exponentially.
@@
We now show that the two contributions to the OTOC decay with the same rate, which coincides with (twice) the leading RP resonance $2\varepsilon$. This feature of the OTOC was observed before for single-particle systems [12], but our results show that it also extends to the many-body setting, suggesting its universality in quantum dynamics. The transfer-matrix expressions in Eqs. (III.4.1) and (62) are exact but their large-$L$, large-$t$ asymptotics are hard to analyze. Nevertheless, we have already shown in the previous section that, in the dissipationless limit the decay rate of the four-point contribution $\overline{\langle\mathcal{O}(y,t)P_{\alpha^{x}}\mathcal{O}(y,t)P_{\alpha^{x}}\rangle}$ is indeed given by $2\varepsilon$.
@@
Figure 5: Exact transfer-matrix result for the decay of the norm. The colored lines give the decay rate $\Delta_{\mathrm{norm}}$ of $\overline{\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle}$ for PBCs in units of $\varepsilon=1$, obtained for $140120<t<$ 140 and $8020<L<$ 80, as a function of $\gamma$. The dashed black line is given by the extrapolation of the collapsed $\Delta_{\mathrm{norm}}(\gamma)$ down to $\gamma\to 0$, where it intercepts the vertical axis at $\Delta_{\mathrm{norm}}(0)=4$.
@@
We thus proceed with the operator norm contribution $\overline{\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle}$, Eq. (62). Exact evaluation of $[\widehat{S}^{L}_{\mathrm{d}}]_{00}$ shows that for $L>2(t-1)$ it becomes independent of $L$ and we have
$\overline{\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle}=(\omega\rho^{2})^{(t-1)}\left[\left(1-\rho^{-1};\omega\right)_{t-1}\right]^{2},$ (64)
@@
where $(a;q)_{n}=\prod_{k=0}^{n-1}(1-aq^{k})$ is the $q$-Pochhammer symbol. Notice that here dissipation acts as a $q$-deformation of the closed-system operator norm. Taking the limit $\omega\to 1$ (i.e., $\gamma=0$) reduces it to the result of the previous section, $\overline{\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle}=1$ regardless of whether the thermodynamic limit is taken before or after, since it is independent of $L$. Instead, we are interested in the long-time decay of the correlator and therefore have to take the limit $t\to\infty$ before $L\to\infty$, which is incompatible with the restriction $L>2(t-1)$ in Eq. (64). In the opposite case, $2(t-1)>L$, we are unable to find a simple closed-form expression. However, we can still evaluate Eq. (62) exactly for arbitrary $t$, $L$, and $\omega$ and extract the decay rate by fitting the late-time decay to an exponential, $\overline{\langle\mathcal{O}(y,t)\mathcal{O}(y,t)\rangle}=A_{\mathrm{norm}}e^{-\Delta_{\mathrm{norm}}t}$, with $A_{\mathrm{norm}}$ and $\Delta_{\mathrm{norm}}$ fitting parameters, as shown in Fig. 5. We find the behavior expected for anomalous relaxation [17]: the curves for different $L$ collapse to a universal curve up to a critical value of the dissipation strength below which the system sees the finite spacing of the lattice and, correspondingly, the gap closes. As $L$ increases (but keeping it below the times used for the fit), this critical point decreases until reaching $p=0$ in the thermodynamic limit. Extrapolating the collapsed curve to this limit, we obtain a decay rate of $\Delta_{\mathrm{norm}}\approx 4\varepsilon$ for $\gamma=0$, as expected.